ORTHO-PARA CONVERSION IN COLLISIONS OF H_{2} WITH H_{2}

\underline{D. L. HUESTIS}, SRI International, Menlo Park, CA 94025.

Molecular hydrogen is the main constituent of the atmospheres of the giant planets, Jupiter, Saturn, Uranus, and Neptune. The relative abundances of ortho- and parahydrogen are inferred from infrared emission spectra of the S(1) and S(0) quadrupole transitions near 17 and 28 μ, sampling conditions in the stratosphere and upper troposphere [1,2]. At most altitudes and latitudes the ortho/para ratio is not in statistical equilibrium at the local temperature, interpreted as resulting from vertical transport from lower hotter or colder regions. Modeling atmospheric circulation requires a quantitative understanding of the rates and mechanisms of ortho-para conversion, which appears to take roughly 30 to 100 years.

The two candidate mechanisms are collisions with paramagnetic aerosols and with the weak magnetic moment of ortho-H_{2}. The better known mechanisms involving ions or H atoms or 4-center atom exchange are inoperative at the relevant low altitudes and low temperatures. An important constraint on atmospheric models could be provided by a quantum mechanical treatment of nuclear spin coupling in collisions of ortho-H_{2} with ortho-H_{2} and para-H_{2}. The only estimate in the literature [1] is based on scaling the magnetic moments of H_{2}(J=1) and O_{2}, which suggests a ortho-para conversion probability of about 4×10^{−19} and a rate coefficient of about 1×10^{−28} cm^{3} s^{−1}.

We will discuss the expected hamiltonian operators involving nuclear spin, procedures for evaluation of matrix elements, and dynamics approaches. Especially interesting are ideas about virtual excitation of the b^{3}Σ_{u}^{+} state resulting in orthopara coupling through nuclear-spin/electron-spin hyperfine interaction [3,4].

[1] B. J. Conrath and P. J. Gierasch, Icarus 57, 184 (1984).

[2] T. Fouchet, E. Lellouch, and H. Feuchtgruber, Icarus 161, 127 (2003).

[3] E. Ilisca and S. Sugano, Phys. Rev. Lett. 57, 2590 (1986).

[4] B. F. Minaev and H. Ågren, J. Phys. Chem. 99, 8396 (1995).